OpenStudy (anonymous):

∠a and ∠b are supplementary. If m∠a = 2x + 1 and m∠b = 3x - 6, what is the measure of ∠b?

OpenStudy (anonymous):

Supplementary would mean that both angles would add up to 180. hence, (2x + 1) + (3x - 6) = 180 2x + 3x = 180 - 5 5x = 175 x = 35 as b is 3x - 6, 35 x 3 = 105. 105 - 6 = 99. b = 99 :)